OpenStudy (anonymous):

Solve by factoring: 3x2 + 2x - 5 = 0. be sure to show your factored expression and then use factors to solve for x Solve using quadratic formula. PLEASE HELP, IT'S TIMED AND I REALLY NEED TO GET IT SOLVED

OpenStudy (carmuz):

Equation have the form: \[ax ^{2} + bx + c = 0\]And solutions are: \[x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\] In your equation: a = 3; b = 2; c = -5 Then: \[ x = \frac{ -2 \pm \sqrt{2^2 - 4(3)(-5)} }{ 2(3) }\]\[ x = \frac{ -2 \pm \sqrt{4 + 60} }{ 2(3) }\] \[x = \frac{ -2 \pm \sqrt{64} }{ 2(3) }\]\[ x = \frac{ -2 \pm 8 }{ 6 }\] Roots: x = {1 ; -5/3}. Factorization: \[\left( x-1 \right)\left( x + \frac{ 5 }{ 3 } \right) = 0\]\[\left( x-1 \right)\left( 3x + 5 \right) = 0\]